# Cyclic Representations of $U_{q}(\widehat{\mathfrak{sl}}_{2})$ and its Borel Subalgebras at Roots of Unity and Q-operators

### v4

@@ 1.1. Background
The key to understanding the algebraic construction and properties of the Q-operator for $U_{q}(\widehat{\mathfrak{sl}}_{2})$ (generated by $e_{i},f_{i},t_{i}^{\pm 1}$, $i\in\{0,1\}$) with $q$ generic is to consider infinite-dimensional representations of the Borel subalgebra $U_{q}(\mathfrak{b}_{+})$ generated by $e_{i},t_{i}^{\pm 1}$ [BLZ96, BLZ97]. Adopting the notation of [CVW24] , there are precisely two independent infinite-dimensional $U_{q}(\mathfrak{b}_{+})$ representations $\rho_{z}$ and $\bar{\rho}_{z}$ that are not restrictions of representations of the full algebra $U_{q}(\widehat{\mathfrak{sl}}_{2})$. Two corresponding Q-operators, ${\mathcal{Q}}(z)$ and ${\widebar{{\mathcal{Q}}}}(z)$ are defined as transfer matrices (i.e. twisted traces over the auxiliary space of products of L-operators) by choosing the auxiliary spaces as $\rho_{z}$ and $\bar{\rho}_{z}$ respectively. In this story, there are also three other representations that may be chosen as auxiliary spaces: $\nu^{\mu}_{z}$ an infinite-dimensional Verma module representation of $U_{q}(\widehat{\mathfrak{sl}}_{2})$ that depends on a complex weight $\mu$; $\pi_{z}$ the two dimensional evaluation representation of $U_{q}(\widehat{\mathfrak{sl}}_{2})$; and $\phi_{z}$ an infinite-dimensional representation of $U_{q}(\mathfrak{b}_{+})$ which is triangular (meaning that $e_{1}$ acts non-trivially, but $e_{0}$ acts as 0). There are two relations between these different representations that give rise to corresponding properties of transfer matrices and Q-operators. They are:
@@
(ii) A short exact sequence    00{0}${\rho_{zq}}$${\rho_{z}\otimes\pi_{z}}$${\rho_{zq^{-1}}}$00{0}$\scriptstyle{\iota}$$\scriptstyle{\tau}$
@@
where $\alpha(z)$ and $\beta(z)$ are rational functions of $z$. Relation (1.1) can be viewed as more fundamental because (1.2) can be derived from it by considering integer $\mu$ (at which points $\nu^{\mu}_{z}$ becomes reducible).
@@ 1.2. Main Results
The algebraic construction of Q-operators for $q^{N}=1$ has been studied in a number of papers in the 2000s, notably in [Kor03a, Kor03b, Kor04, Roa06b, Roa06a]. However, an understanding of the analogues of property (i) and Equation (1.1) in terms of representations of the Borel subalgebra is lacking. This paper seeks to fill this gap.
@@
This is the analog of property (i) above. We then go on in Theorem 3.7 to give $U_{q}(\mathfrak{b}_{+})$ short exact sequences that are analogous to (ii) above. Parallel results that give these factorization and fusion properties in terms of L-operators are then given in Propositions 3.13 and 3.14 respectively. In Section 4 we use these various representations as auxiliary spaces in the definition of Q operators and transfer matrices.
@@ 2.1. Cyclic Representations and R-matrices of $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$
We closely follow the notation and conventions of Section 4 of [DJMM91b]. We begin by defining $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$, which is an algebra over $\mathbb{C}$ generated by $e_{i},f_{i},t_{i}^{\pm 1},z_{i}$ $(i=0,1)$, where $e_{i},f_{i},t_{i}^{\pm 1}$ satisfy the standard relations of the quantum affine algebra $U_{q}(\widehat{\mathfrak{sl}}_{2})$, and $z_{0}$ and $z_{1}$ are two new central elements. The comultiplication of $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ is chosen as
@@ 2.1.1. Cyclic Representations $\Omega_{rs}$
where $k$ is the modulus with $k^{2}+k^{\prime 2}=1$. With the notation $r=(x_{r},y_{r},\mu_{r})$ and $s=(x_{s},y_{s},\mu_{s})$, the representation $\Omega_{rs}:\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})\rightarrow\hbox{End}(W)$ is given by 111 To avoid confusion, we note that our $\Omega_{rs}$ corresponds to $\xi=sr$ in the notation of [DJMM91b]. The other difference in notation is that $Z$ of [DJMM91b] is our $Z^{2}$ the reason for this change will become apparent when we consider L-operators in Section 3.3.
@@ 2.1.2. R-Matrices $\check{R}(rr^{\prime};ss^{\prime})$
The $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ representation $\Omega_{rs}$ was constructed, and parameterized in terms of $(r,s)\in{\mathcal{C}}_{k}\times{\mathcal{C}}_{k}$, precisely in order that an invertible $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ intertwiner
@@
Of course, the pair of properties 2.4 and 2.5 is the same as the statement that there exists a $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ isomorphism $\check{R}(rr^{\prime};ss^{\prime}):\Omega_{rr^{\prime}}\otimes\Omega_{ss^{\prime}}\rightarrow\Omega_{ss^{\prime}}\otimes\Omega_{rr^{\prime}}$.
@@
Proposition 2.1 (DJMMb).
@@
The full proof is given in [DJMM91b]. Here, we note that assuming $T_{rs}$ and $S_{rs}$ are of the form (2.7) then it follows that they are $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ intertwiners of the form stated provided they satisfy respectively
@@
The coefficients $\overline{W}_{rs}(n)$ and $W_{rs}(n)$ are the Boltzmann weights of the chiral Potts model [BPAY88], and Proposition 2.1 is the algebraic statement of the factorization property of the R-matrix first observed in [BS90]. The discrete Fourier transform $\widecheck{W}_{rs}(n)$ will be used in the proof in Appendix A.3. We fix the normalization of our weights by the choice
@@
$r^{\prime}$$r\,$$s^{\prime}$$r^{\prime}$$s$$s^{\prime}$$r$$s$$S_{r^{\prime}s}\sim$$:\Omega_{rr^{\prime}}\otimes\Omega_{ss^{\prime}}\rightarrow\Omega_{rs}\otimes\Omega_{r^{\prime}s^{\prime}}$
$s$$r$$r$$s$$T_{rs}\sim$$:\Omega_{rs}\rightarrow\Omega_{sr}$
@@
$\check{B}_{r^{\prime};ss^{\prime}}=(\mathbb{I}\otimes T_{r^{\prime}s^{\prime}})S_{r^{\prime}s}=$$:\Omega_{rr^{\prime}}\otimes\Omega_{ss^{\prime}}\rightarrow\Omega_{rs}\otimes\Omega_{s^{\prime}r^{\prime}}$$r\,$$r^{\prime}$$s$$s^{\prime}$$r$$s$$s^{\prime}$$r^{\prime}$
$r^{\prime}$$r\,$$s^{\prime}$$r$$s$$s^{\prime}$$r^{\prime}$$s$$\check{A}_{rr^{\prime};s}=S_{rs}(T_{rr^{\prime}}\otimes\mathbb{I})=$$:\Omega_{rr^{\prime}}\otimes\Omega_{ss^{\prime}}\rightarrow\Omega_{r^{\prime}s}\otimes\Omega_{rs^{\prime}}$

### v5

@@ 1.1. Background
The key to understanding the algebraic construction and properties of the Q-operator for $U_{q}(\widehat{\mathfrak{sl}}_{2})$ (generated by $e_{i},f_{i},t_{i}^{\pm 1}$, $i\in\{0,1\}$) with $q$ generic is to consider infinite-dimensional representations of the Borel subalgebra $U_{q}(\mathfrak{b}_{+})$ generated by $e_{i},t_{i}^{\pm 1}$ [BLZ96, BLZ97]. Adopting the notation of [CVW24] , there are precisely two independent infinite-dimensional $U_{q}(\mathfrak{b}_{+})$ representations $\rho_{z}$ and $\bar{\rho}_{z}$ that are not restrictions of representations of the full algebra $U_{q}(\widehat{\mathfrak{sl}}_{2})$ [^1] The representations can be extended to the algebras known as the asymptotic algebra and shifted quantum affine algebra - see [HJ12] and [H20]. . Two corresponding Q-operators, ${\mathcal{Q}}(z)$ and ${\widebar{{\mathcal{Q}}}}(z)$ are defined as transfer matrices (i.e. twisted traces over the auxiliary space of products of L-operators) by choosing the auxiliary spaces as $\rho_{z}$ and $\bar{\rho}_{z}$ respectively. Throughout this work we restrict to $N$ odd and $N\geq 3$; the representation theory for $N$ odd and $N$ even has significant differences as explained in and [DCK92] and [Kor03a]. In our story, there are also three other representations that may be chosen as auxiliary spaces: $\nu^{\mu}_{z}$ - an infinite-dimensional Verma module representation of $U_{q}(\widehat{\mathfrak{sl}}_{2})$ that depends on a complex weight $\mu$; $\pi_{z}$ - the two dimensional evaluation representation of $U_{q}(\widehat{\mathfrak{sl}}_{2})$; and $\phi_{z}$ - an infinite-dimensional representation of $U_{q}(\mathfrak{b}_{+})$ which is triangular (meaning that $e_{1}$ acts non-trivially, but $e_{0}$ acts as 0). There are two relations between these different representations that give rise to corresponding properties of transfer matrices and Q-operators. They are:
@@
[…]
@@
where $\alpha(z)$ and $\beta(z)$ are rational functions of $z$. A similar relation to (1.2) is satisfied by ${\widebar{{\mathcal{Q}}}}(z)$. Relation (1.1) can be viewed as more fundamental because (1.2) can be derived from it by considering integer $\mu$ (at which points $\nu^{\mu}_{z}$ becomes reducible).
@@ 1.2. Main Results
The algebraic construction of Q-operators for $q^{N}=1$ has been studied in a number of papers in the 2000s, notably in [Kor03a, Kor03b, Kor04a, Roa06b, Roa06a]. One of the key motivations was to understand the degeneracies in the spectrum of the $XXZ$ Hamiltonian at roots of unity observed in [DFM01]. While an algebraic understanding of the roots of unity analogue of (1.2 ) was obtained in this earlier body of work [Kor03a], analogues of property (i) and Equation (1.1) in terms of representations of the Borel subalgebra were lacking. This paper seeks to fill this gap.
@@
This is the analog of property (i) above. We then go on in Theorem 3.8 to give $U_{q}(\mathfrak{b}_{+})$ short exact sequences that are analogous to (ii) above. Parallel results that give these factorization and fusion properties in terms of L-operators are then given in Propositions 3.14 and 3.15 respectively. In Section 4 we use these various representations as auxiliary spaces in the definition of Q operators and transfer matrices.
@@ 1.3. Goals and Motivation
1.4. Summary of Notation
Notation Description Diagrammatics Section
$q$ primitive $N$’th root of unity, $N$ odd $\geq 3$
$\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ central extension of quantum affine algebra $U_{q}(\widehat{\mathfrak{sl}}_{2})$ 2.1
$U_{q}(\mathfrak{b}_{+})$ Borel sualgebra of $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ 2.2
$W$, $V$ $W=\mathbb{C}^{N}$, $V=\mathbb{C}^{2}$ 2.1.1
$X,Z$ elements of $GL(W)$ satisfying Eqn. (2.1) 2.1.1
${\mathcal{C}}_{k}$ chiral Potts algebraic curve 2.1.1
$\Omega_{rs}$ cyclic repn: $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})\rightarrow\operatorname{End}(W)$; $\in{\mathcal{C}}_{k} ss\,rr\,$ 2.1.1
$\rho_{r}$ cyclic repn: $U_{q}(\mathfrak{b}_{+})\rightarrow\operatorname{End}(W)$; $\in{\mathcal{C}}_{k} rr\,$ 2.2
$\bar{\rho}_{r}$ cyclic repn: $U_{q}(\mathfrak{b}_{+})\rightarrow\operatorname{End}(W)$; $\in{\mathcal{C}}_{k} rr\,$ 2.2
$\varphi_{c}$ repn: $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})\rightarrow\operatorname{End}(W)$ (dir. sum of 1D repns); $\in\mathbb{C}^{\times} cc\,$ 2.2
$\pi_{z}$ 2D evaluation repn: $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})\rightarrow\operatorname{End}(\mathbb{C}^{2})$; $\in\mathbb{C}^{\times} zz\,$ 2.2
$T_{rs}$ $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ int: $\Omega_{rs}\rightarrow\Omega_{sr}$ ssrrrrss 2.1.2
$S_{rs}$ $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ int: $\Omega_{rr^{\prime}}\otimes\Omega_{ss^{\prime}}\rightarrow\Omega_{rs}\otimes\Omega_{r^{\prime}s^{\prime}}$ $^{\prime}rr\,$$s^{\prime}$$r^{\prime}$$ss^{\prime}$rrss 2.1.2
$\check{B}_{r^{\prime};ss^{\prime}}$ $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ int $(\mathbb{I}\otimes T_{r^{\prime}s^{\prime}})S_{r^{\prime}s}:\Omega_{rr^{\prime}}\otimes\Omega_{ss^{\prime}}\rightarrow\Omega_{rs}\otimes\Omega_{s^{\prime}r^{\prime}}$ $rr^{\prime}$$ss^{\prime}$$ss^{\prime}rr^{\prime}$ 2.1.2
$\check{A}_{rr^{\prime};s}$ $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ int $S_{rs}(T_{rr^{\prime}}\otimes\mathbb{I}):\Omega_{rr^{\prime}}\otimes\Omega_{ss^{\prime}}\rightarrow.\Omega_{r^{\prime}s}\otimes\Omega_{rs^{\prime}}$ $^{\prime}rr\,$$s^{\prime}$$s^{\prime}rr^{\prime}ss$ 2.1.2
$\check{R}(rr^{\prime};ss^{\prime})$ $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ int $S_{rs^{\prime}}(T_{rs}\otimes T_{r^{\prime}s^{\prime}})S_{r^{\prime}s}:\Omega_{rr^{\prime}}\otimes\Omega_{ss^{\prime}}\rightarrow\Omega_{ss^{\prime}}\otimes\Omega_{rr^{\prime}}$ see text 2.1.2
$\overline{W}_{rs}(n),\,\widehat{W}_{rs}(n)$ Components of $T_{rs}$ and $S_{rs}$ = Chiral Potts Boltz. wts 2.1.2
$W_{rs}(n)$, $\widecheck{W}_{rs}(n)$ Transformed versions of $\overline{W}_{rs}(n)$, $\widehat{W}_{rs}(n)$ 2.1.2
$\chi$ $X^{-1}\otimes X$ 3.1
${\mathcal{O}}(\chi)$ $U_{q}(\mathfrak{b}_{+})$ int $\Omega_{rs}\otimes\varphi_{c_{0}}\rightarrow\rho_{r}\otimes\bar{\rho}_{s}$ rrssrrss 3.1
${\mathcal{P}}(Z)$ polynomial $\sum\limits_{n=0}^{N-1}p_{n}Z^{2n}$ 3.1
$\mathfrak{T}_{rs}$ $U_{q}(\mathfrak{b}_{+})$ int $\rho_{r}\otimes\bar{\rho}_{s}\rightarrow\rho_{s}\otimes\bar{\rho}_{r},$ 3.1
${\mathfrak{S}}_{rs}$ $U_{q}(\mathfrak{b}_{+})$ int $\bar{\rho}_{r}\otimes\rho_{s}\rightarrow\bar{\rho}_{s}\otimes\rho_{r}.$ 3.1
Notation Description Diagrammatics Section
[…]
$\bar{\iota}_{s}$ $U_{q}(\mathfrak{b}_{+})$ int $\bar{\rho}_{sq}\rightarrow\bar{\rho}_{s}\otimes\pi_{z_{s}}$ see text 3.2
$\bar{I}_{s}$ $U_{q}(\mathfrak{b}_{+})$ int $\Omega_{r,sq}\rightarrow\Omega_{r,s}\otimes\pi_{z_{s}}$ see text 3.2
[…]
$\bar{\tau}_{s}$ $U_{q}(\mathfrak{b}_{+})$ int $\bar{\rho}_{s}\otimes\pi_{z_{s}}\rightarrow\bar{\rho}_{sq^{-1}}$ see text 3.2
$\bar{T}_{s}$ $U_{q}(\mathfrak{b}_{+})$ int $\Omega_{r,s}\otimes\pi_{z_{s}}\rightarrow\Omega_{r,sq^{-1}}$ see text 3.2
$\check{L}_{\Omega_{rs}}(z)$ $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ int $\Omega_{rs}\otimes\pi_{z}\rightarrow\pi_{z}\otimes\Omega_{rs}$ ss\,rr\,$z$ 3.3
$\check{L}_{\varphi}$ $U_{q}(\mathfrak{b}_{+})$ int $\varphi_{c}\otimes\pi_{z}\rightarrow\pi_{z}\otimes\varphi_{c}$ $z$ 3.3
$\check{L}_{\rho_{r}}(z)$ $U_{q}(\mathfrak{b}_{+})$ int $\rho_{r}\otimes\pi_{z}\rightarrow\pi_{z}\otimes\rho_{r}$ rr\,$z$ 3.3
$\check{L}_{\bar{\rho}_{r}}(z)$ $U_{q}(\mathfrak{b}_{+})$ int $\bar{\rho}_{r}\otimes\pi_{z}\rightarrow\pi_{z}\otimes\bar{\rho}_{r}$ rr\,$z$ 3.3
$\check{\bf L}_{\Omega_{rs}}(z)$ $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ int $\pi_{z}\otimes\Omega_{rs}\rightarrow\Omega_{rs}\otimes\pi_{z}$ zz\,rr\,$s$ 3.3
$\check{R}(z/w)$ $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ int $\pi_{z}\otimes\pi_{w}\rightarrow\pi_{w}\otimes\pi_{z}$ ; 6-vertex model R-matrix zzww 3.3
$U_{r}(z)$, $V_{r}(z)$ $2\times 2$ matrices; $r\in{\mathcal{C}}_{k}$, $z\in\mathbb{C}^{\times}$ 3.3
$\{A,B\}$ Element of $\operatorname{End}(W\otimes V)$ defined by Eqn (3.7) 3.3
$T(z)$ $\in\operatorname{End}(V^{\otimes M})$ ; 6-vertex model transfer matrix see text 4.1
$Q_{\rho_{r}}(w)$ $\in\operatorname{End}(V^{\otimes M})$ ; 6-vertex model Q-operator see text 4.1
$Q(z,\mu)$ transformed version of $Q_{\rho_{r}}(w)$ see text 4.1
$Q_{\bar{\rho}_{r}}(w)$ $\in\operatorname{End}(V^{\otimes M})$ ; 6-vertex model Q-operator see text 4.1
$\widebar{Q}(z,\mu)$ transformed version of $Q_{\bar{\rho}_{r}}(w)$ see text 4.1
$T_{\Omega_{rs}}(w)$ $\in\operatorname{End}(V^{\otimes M})$ see text 4.1
$T_{\varphi_{c}}(w)$ $\in\operatorname{End}(V^{\otimes M})$ see text 4.1
${\mathcal{T}}(z)$ $\in\operatorname{End}(W^{\otimes M})$ ; $\tau_{2}$ model transfer matrix see text 4.2
${{\mathcal{Q}}}_{r^{\prime};ss^{\prime}}$ $\in\operatorname{End}(W^{\otimes M})$ ; $\tau_{2}$ model Q-operator see text 4.2
${\mathcal{T}}^{CP}$ $\in\operatorname{End}(W^{\otimes M})$ ; chiral Potts model transfer matrix see text 4.2
@@ 2.1. Cyclic Representations and R-matrices of $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$
We closely follow the notation and conventions of Section 4 of [DJMM91b] who in turn developed the earlier work of [DCK92] on representations of quantum groups at roots of unity. We make use the algebra $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ defined in Section 4 of [DJMM91b]: it is an algebra over $\mathbb{C}$ generated by $e_{i},f_{i},t_{i}^{\pm 1},z_{i}$ $(i=0,1)$, where $e_{i},f_{i},t_{i}^{\pm 1}$ satisfy the standard relations of the quantum affine algebra $U_{q}(\widehat{\mathfrak{sl}}_{2})$, and $z_{0}$ and $z_{1}$ are two new central elements. The comultiplication of $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ is chosen as
@@ 2.1.1. Cyclic Representations $\Omega_{rs}$
where $k$ is the modulus with $k^{2}+k^{\prime 2}=1$. With the notation $r=(x_{r},y_{r},\mu_{r})$ and $s=(x_{s},y_{s},\mu_{s})$, the representation $\Omega_{rs}:\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})\rightarrow\hbox{End}(W)$ is given by  [^2] To avoid confusion, we note that our $\Omega_{rs}$ corresponds to $\xi=sr$ in the notation of [DJMM91b]. The other difference in notation is that $Z$ of [DJMM91b] is our $Z^{2}$ - the reason for this change will become apparent when we consider L-operators in Section 3.3.
@@ 2.1.2. R-Matrices $\check{R}(rr^{\prime};ss^{\prime})$
The $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ representation $\Omega_{rs}$ was constructed, and parameterized in terms of $(r,s)\in{\mathcal{C}}_{k}\times{\mathcal{C}}_{k}$, precisely in order that an invertible $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ linear map
@@
As we shall be discussing many such intertwiners in this paper, it is useful at this point to introduce some abbreviated notation: the pair of properties (2.4) and (2.5) is of course the same as the statement that there exists a isomorphism of $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ representations
(2.6) $\check{R}(rr^{\prime};ss^{\prime}):(\Omega_{rr^{\prime}},W)\otimes(\Omega_{ss^{\prime}},W)\rightarrow(\Omega_{ss^{\prime}},W)\otimes(\Omega_{rr^{\prime}},W),$
where the representation $(\Omega_{rs},W)$ consists of the vector space $W$ and the algebra action $\Omega_{rs}:\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})\rightarrow\operatorname{End}(W)$. Our abbreviation is to express relation (2.6) as
(2.7) $\check{R}(rr^{\prime};ss^{\prime}):\Omega_{rr^{\prime}}\otimes\Omega_{ss^{\prime}}\rightarrow\Omega_{ss^{\prime}}\otimes\Omega_{rr^{\prime}}.$
@@
Proposition 2.1 ( Date, Jimbo, Miki, Miwa).
@@
This statement is Theorem 4.2 of [DJMM91b] and the full proof can be found there. Here, we note that assuming $T_{rs}$ and $S_{rs}$ are of the form (2.9) then it follows that they are $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ intertwiners of the form stated provided they satisfy respectively
@@
The coefficients $\overline{W}_{rs}(n)$ and $W_{rs}(n)$ are the Boltzmann weights of the chiral Potts model [BPAY88], and Proposition 2.1 is the algebraic statement of the factorization property of the R-matrix first observed in [BS90].  [^3] The additional central elements $z_{0},z_{1}$ were included in the definition of $\widetilde{U}_{q}(\widehat{\mathfrak{sl}}_{2})$ in [DJMM91b] in order to achieve this factorization into the standard Chiral Potts weights. The discrete Fourier transform $\widecheck{W}_{rs}(n)$ will be used in the proof in Appendix A.3. We fix the normalization of our weights by the choice
@@
$^{\prime}rr\,$$s^{\prime}$$r^{\prime}$$ss^{\prime}$$rsS_{r^{\prime}s}\sim$$\Omega_{rr^{\prime}}\otimes\Omega_{ss^{\prime}}\rightarrow\Omega_{rs}\otimes\Omega_{r^{\prime}s^{\prime}}$
[…]
@@
$\check{B}_{r^{\prime};ss^{\prime}}=(\mathbb{I}\otimes T_{r^{\prime}s^{\prime}})S_{r^{\prime}s}=$$\Omega_{rr^{\prime}}\otimes\Omega_{ss^{\prime}}\rightarrow\Omega_{rs}\otimes\Omega_{s^{\prime}r^{\prime}}$$rr^{\prime}$$ss^{\prime}$$ss^{\prime}rr^{\prime}$
$^{\prime}rr\,$$s^{\prime}$$s^{\prime}rr^{\prime}ss$$\check{A}_{rr^{\prime};s}=S_{rs}(T_{rr^{\prime}}\otimes\mathbb{I})=$$\Omega_{rr^{\prime}}\otimes\Omega_{ss^{\prime}}\rightarrow\Omega_{r^{\prime}s}\otimes\Omega_{rs^{\prime}}$
@@
(2.18) $^{\prime}rr\,$$^{\prime}rr$$^{\prime}ssss^{\prime}$$\check{R}({rr^{\prime};ss^{\prime}})=\check{A}_{rs;s^{\prime}r^{\prime}}\circ\check{B}_{r^{\prime};ss^{\prime}}=$$\Omega_{rr^{\prime}}\otimes\Omega_{ss^{\prime}}\rightarrow\Omega_{ss^{\prime}}\otimes\Omega_{rr^{\prime}}$